OpenStudy (anonymous):

x^3-x^2-36x+36/x^3-17x^2+90x-144 whats a vertical asymtome and the point were the hole is

OpenStudy (anonymous):

holes=3,6,8

OpenStudy (anonymous):

it says has one hole at the point:(______)(______)

OpenStudy (anonymous):

it has two wertical asymtote.. x=3 and x=8

OpenStudy (anonymous):

how did you get the 3 if u can explain i got a 6 and a 8 but the 6 was wrong

OpenStudy (anonymous):

\[\frac{x^3-x^2-36x+36}{x^3-17x^2+90x-144}=\frac{(x-6)(x-1)(x+6)}{(x-8)(x-6)(x-3)}\]

OpenStudy (anonymous):

x-6 cancel out..

OpenStudy (anonymous):

oh i see thanks

OpenStudy (anonymous):

np